OpenStudy (wolframwizard):

Find the value(s) of k such that the graphs of the quadratic functions f(x)=(x+3)^2+k and g(x)=-(x+k)^2+3 are tangent to one another.

OpenStudy (solomonzelman):

is this calculus?

OpenStudy (wolframwizard):

No this should be solved using Algebra 2

OpenStudy (wolframwizard):

I've tried making the equations equal to each other and solving for k by setting the discriminant equal to 0 but my answers were wrong. The correct answer is 1 and 3 but I need to know how to get there

OpenStudy (solomonzelman):

Press play and you will see those answers. https://www.desmos.com/calculator/alw9wn9kk2 I am not really sure how to do this without calculus (i'll look into the problem)

OpenStudy (solomonzelman):

I know how they arrive at k=3 with algebra

OpenStudy (wolframwizard):

3 is easy but I don't know how to get 1

OpenStudy (solomonzelman):

So, you know how to get k=3?

OpenStudy (wolframwizard):

Yes

OpenStudy (solomonzelman):

\(\color{#000000 }{ \displaystyle y=(x+3)^2 }\) \(\color{#000000 }{ \displaystyle y=-(x+k)^2 }\) They are tangent, at k=3, because then they are reflected about x=0, and we know that in that case k=3, so if you say \(\color{#000000 }{ \displaystyle y=(x+3)^2 +k }\) \(\color{#000000 }{ \displaystyle y=-(x+k)^2 +3 }\) then we use the same fact, but this time they are reflected about x=3.

OpenStudy (anonymous):

i think i have seen a problem like this before, but i kind of forget how to do it i think it has something to do with setting them equal, then making sure they intersect at only one point by making sure the discriminant is zero

OpenStudy (anonymous):

we can try it and see if we get the two answers you know to be true

OpenStudy (wolframwizard):

I tried that but the equation for k didn't give correct values. I'll try again though

OpenStudy (anonymous):

did you start with \[-(x+k)^2+3=(x+3)^2+k\]?

OpenStudy (anonymous):

i need paper to do this, hold the phone

OpenStudy (wolframwizard):

I tried it again and got the correct answers. I must've just done some calculation error on the first attempt sorry

OpenStudy (anonymous):

i did too, let me try again

OpenStudy (anonymous):

ok not the second time either damn

OpenStudy (anonymous):

third time is a charm! got it

OpenStudy (anonymous):

start here \[-(x+k)^2+3=(x+3)^2+k\]

OpenStudy (anonymous):

expand and set equal to zero you want my work?

OpenStudy (anonymous):

@SolomonZelman i have seen this before, to do without calc think i googled it or something ideas is to set them equal, then take the discriminant, set it equal to zero and solve, thereby making sure there is only one solution to the intersection i.e. they are tangent

OpenStudy (anonymous):

confused the hell out of me the first time i saw it

OpenStudy (anonymous):

@WolframWizard you get it yet or no?

OpenStudy (solomonzelman):

Yes, satelitte, when you said this thing about setting the discriminant zero I figured it out ... just that not everytime I can come up with something; because figuring after you've seen someone is easier. tnx for the information!

OpenStudy (anonymous):

turns out in this case if you set \(b^2-4ac=0\) you get \(k=3,k=1\) never thought of your other reason before wonder if you can come up with the 1 by sheer reason as well?

OpenStudy (solomonzelman):

well I wil expect that calculus is sheer.

OpenStudy (anonymous):

lol

OpenStudy (solomonzelman):

But, not for this post since the user needs to solve with algebra...

OpenStudy (solomonzelman):

perhaps there is some other way of solving it besides calc and besides \(\Delta =0\)